Pearl Inc. owns the following assets. Estimated Useful Life salvager 10 years 5 years 12 years Asset Cost $72,200 $7,220 53,300 ,330 C 100,0404,880 A Compute the composite depreciation rate and the composite life of Pearls assets. (Round answers to i Composite depreciation rate Composite life years
Answer #1

Please find the detailed answer as follows:

Depreciation Expense (A) = (72200 - 7220)/10 = 6,498

Depreciation Expense (B) = (53300 - 5330)/5 = 9,594

Depreciation Expense (C) = (100,040 - 4880)/12 = 7,930

Total Depreciation Expense = 6498 + 9594 + 7930 = 24,022

Composite depreciation rate = 24022/(72,200 + 53300 + 100,040)*100 = 10.65%

-----

Composite life = (72,200 - 7220 + 53,300 - 5330 + 100,040 - 4880)/24,022 = 8.66 years